OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of square root of the quantity 16 r to the ninth power. The total number of tiles used can be represented by the expression below. one hundred twenty r to the third power, all over the square root of the quantity sixteen r to the ninth power

OpenStudy (studygurl14):

\[120r^{3} = \sqrt{16r^{9}}\] Is that the equation?

OpenStudy (anonymous):

\[\frac{ 120r^2}{ \sqrt{16r^9}}\]

OpenStudy (studygurl14):

oh! okay \[\frac{ 120r^{3} }{ \sqrt{16r^{9}} } = \frac{ 120r^{3} }{ 16^{1/2}r^{9/2} }\] Wait, is it 120r^2, or 120r^3

OpenStudy (anonymous):

Its 120^2

OpenStudy (studygurl14):

okay, \[\frac{ 120r^{2} }{ \sqrt{16r^{9}} } = \frac{ 120r^{2} }{4r^\frac{ 9 }{ 2 } }\]

OpenStudy (anonymous):

I don't know what you did here

OpenStudy (studygurl14):

\[\sqrt{16} = 4\] right? Well i just took that out also... \[\sqrt[n]{x^{m}} = x^{\frac{ m }{ n }}\] and since... \[\sqrt{x} = \sqrt[2]{x}\] \[\sqrt{r^{9}} = r^{\frac{ 9 }{ 2 }}\] see?

OpenStudy (anonymous):

Okay... I got 30... But when I look around I see 7.5. Which is right?

OpenStudy (studygurl14):

it's 30 but don't forget about the r

OpenStudy (anonymous):

Don't the r's cancel out?

OpenStudy (anonymous):

@april115 Can you help with this? Please?

OpenStudy (anonymous):

@Agent47 Help?

OpenStudy (anonymous):

@adrynicoleb Can you PLEASEE help!?

OpenStudy (adrynicoleb):

Ok. I'll see what I can do.

OpenStudy (anonymous):

Thank you SOO much!

OpenStudy (adrynicoleb):

But I'm warning you, I am not that good at math. ._.

OpenStudy (anonymous):

I'm not either. (:

OpenStudy (adrynicoleb):

Lol yeah. I don't know. *bows head in shame*

OpenStudy (anonymous):

.... Uh. This is so hard!

OpenStudy (studygurl14):

\[\frac{ 120r^{2} }{ 4r^{9/2} } = \frac{ 30r^{2} }{ r^{9/2} } = \frac{ 30r^{2} }{ \sqrt{r^{9}} }\] \[\frac{ 30r^{2} }{ \sqrt{r^{9}}} * \frac{ \sqrt{r^{9}} }{\sqrt{r^{9}} } = \frac{ 30r^{2}\sqrt{r^{9}} }{ r^{9} }\] \[= 30^{-7}\sqrt{r^9}\]